OpenStudy (anonymous):

Please help! only answer if you know what you're talking about. I cannot get this question wrong. A jar contains jelly beans that are red, yellow, blue, and purple. Several people reach in and grab handfuls of jelly beans, then collectively tally the number of each color: 15 red, 20 yellow, 12 blue, and 8 purple. If someone pulls out 22 jelly beans, about how many would you expect to be yellow? A. 20 B. 8 C. 4 D. 13

OpenStudy (anonymous):

Hey I can help

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[15+20+12+8=55\] i believe solve\[\frac{20}{55}=\frac{x}{22}\] for \(x\)

OpenStudy (anonymous):

Is it b or a?

OpenStudy (anonymous):

beware of answers with no explanation they are frequently wrong, and in any case you will have no idea if it is right or not

OpenStudy (anonymous):

Because if there are 20 yellow beans and 22 are pulled out, 20 jelly beans is possible

OpenStudy (anonymous):

Thanks, the answer was 8.

OpenStudy (anonymous):

see what i mean? no clew solve the ratio, you will get the answer

OpenStudy (anonymous):

Clue..

OpenStudy (anonymous):

Well its possible that you can get twenty. In my opinion it could be worded better...

OpenStudy (anonymous):

Actually the answer was 8.

OpenStudy (anonymous):

I know... Im sorry... Do you go to LSS?

OpenStudy (anonymous):

You gave me a wrong answer on purpose?

OpenStudy (anonymous):

No I didn't I thought it was 20 because they didn't say anything about having to be by ratio

OpenStudy (anonymous):

ratios is the best way to do it

OpenStudy (anonymous):

Yeah. Again I'm sorry